Aqueous hydrochloric acid HCl will react with solid sodium hydroxide NaOH

 to produce aqueous sodium chloride NaCl and liquid water H2O. Suppose 8.0 g of hydrochloric acid is mixed with 17.1 g of sodium hydroxide. Calculate the minimum mass of hydrochloric acid that could be left over by the chemical reaction. Round your answer to 2 significant digits. 


1
Expert's answer
2021-12-06T02:45:03-0500

NaOH (aq) + HCl (aq) → NaCl (aq) + H2O (l)



Moles of HCl "= \frac{8}{36}=0.22moles"


Moles of NaOH"=\frac{17.1}{40}=0.43moles"


"0.43-0.22= 0.21moles"


"0.21\u00d7 36=7.7grams"